OpenStudy (anonymous):

using quadratic formula... A car traveled 60 miles at a certain speed. Then is traveled 50 miles at a speed 10 MPH less. If the total time for the trip was 2 hours what was the speed of the first part of the trip?

OpenStudy (anonymous):

how do you write the equation?

OpenStudy (ddcamp):

For each part of the trip, \[time = \frac{distance}{speed}\] You know the distance and speed for the second part. I don't think this actually needs the quadratic formula though.

OpenStudy (anonymous):

ok will try. the problem is that the teacher wants him using quadratic formula, that is the section they are on

OpenStudy (ddcamp):

Hmm... Just realized I misread the problem :P sorry.

OpenStudy (anonymous):

was about to ask.... :)

OpenStudy (ddcamp):

Try setting it up something like: \[2 hours = \frac{60 miles}{x\ mph} + \frac{50miles}{x-1\ mph}\]

OpenStudy (ddcamp):

Where x is the speed on the first part of the trip.

OpenStudy (anonymous):

bottom of the 50 mile side... should that be x-10MPH?

OpenStudy (ddcamp):

Yeah...

OpenStudy (ddcamp):

I should probably get some rest before trying to help anyone else.

OpenStudy (anonymous):

we had that equation at one point but failed to get the answer

OpenStudy (ddcamp):

Try multiplying both sides of the equation by (x-10) and x, to get rid of the denominators. Then, combine like terms.

OpenStudy (anonymous):

so \[60(x-10)+50x=2x(x-10)\]

OpenStudy (anonymous):

we got 60 can anyone check for me?